40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 8095     Entry time: Sat Feb 16 19:23:17 2013
Author: yuta 
Type: Update 
Category: LSC 
Subject: PR2 flipped PRMI locked 

It is my pleasure to announce that the first lock of PR2 flipped PRMI was succeeded.



POP looks very nice. TEM00 and not wobbling.
We need more I/Q phase and gain/filter adjustment and characterization soon.

Some more details:
  MICH error signal: AS55_Q_ERR (using POP55 PD; phase rotation angle 70 deg)
  PRCL error signal: REFL11_I_ERR (phase rotation angle 80 deg)
  MICH feedback: BS (MICH_GAIN = -60)
  PRCL feedback: PRM (PRCL_GAIN = -0.5)

ELOG V3.1.3-